# Cool-Core Destruction in Merging Clusters with AGN Feedback and Radiative Cooling

### v1

@@ 1 Introduction
Over the past decade, simulations of CC clusters incorporating mechanical AGN feedback and radiative cooling have successfully produced self-regulated systems and CC profiles consistently with observations (Gaspari et al. 2011, 2012, Li & Bryan 2014, Prasad et al. 2015, Yang & Reynolds 2016). On the other hand, isolated cluster simulations investigating whether AGN feedback is capable of destroying the CCs have not reached a definitive conclusion. Guo & Oh (2009) and Barai et al. (2016) find that AGN feedback can transform CCs into NCCs, while Gaspari et al. (2013) and Ehlert et al. (2022) show that an efficient self-regulation preserves the CC structure over the long term, in particular through chaotic cold accretion. In order to understand the role of AGN feedback in the CC/NCC dichotomy, it is critical to investigate the interplay between radiative cooling and AGN feedback/feeding during diverse cluster merger events.
@@
The structure of the paper is as follows. In Section 2, we describe the setups of our cluster merger simulations and the modules of AGN feedback and radiative cooling. In Section 3, we present the comparison of merger simulations with different physical processes in Section 3.1, the entropy evolution of mergers with AGN feedback and radiative cooling in Section 3.2, and the processes of heating and cooling in Section 3.3. In Section 4, we compare our results with previous works in Section 4.1, discuss the implications of CC/NCC transitions in Section 4.2, and address the limitations of our simulations in Section 4.3. Finally, we summarize our findings in Section 5.
2 Methods
We perform 3D hydrodynamic simulations using GAMER (v2) (Schive et al., 2018), a GPU-accelerated adaptive mesh refinement (AMR) code. With GPU-acceleration, GAMER outperforms FLASH by nearly two orders of magnitude in cluster merger simulations. Our main focus is the mergers of galaxy clusters incorporating AGN feedback and radiative cooling, but we also conduct mergers only and mergers without AGN feedback for comparison. We vary the mass ratios and impact parameters of cluster mergers to assess their respective impacts on CC destruction.
2.1 Simulation Setup
@@
where $\Psi=-\Phi$ is the relative potential and ${\cal E}=\Psi-\frac{1}{2}v^{2}$ is the relative energy of the particle. We tabulate the function ${\cal F}$ in intervals of ${\cal E}$ interpolate to solve for the distribution function at a given energy. Given the radius of the particle, particle speeds can then be chosen from this distribution function using the acceptance-rejection method. Once particle radii and speeds are determined, positions and velocities are determined by choosing random unit vectors in $\Re^{3}$.
@@
$M_{\mathrm{DM},200}$ ($M_{\odot}$) 5.497×$\times 10^{14}$ 1.872×$\times 10^{14}$ 5.726×$\times 10^{13}$
$M_{\mathrm{gas},200}$ ($M_{\odot}$) 7.786×$\times 10^{13}$ 2.189×$\times 10^{13}$ 5.631×$\times 10^{12}$
$M_{\mathrm{tot},200}$ ($M_{\odot}$) 6.275×$\times 10^{14}$ 2.091×$\times 10^{14}$ 6.289×$\times 10^{13}$
@@
We perform 3D simulations of mergers between a primary and a secondary cluster within a cubic domain of (15 Mpc)3. Diode boundary conditions are adopted to permit outflow while preventing inflow. The cluster mergers are assumed to occur on the $x-y$ plane. For each simulation, the $x$ coordinates of the two merging clusters are fixed at 5.74 Mpc and 9.26 Mpc, respectively, measured from the left edge of the simulation box. The $y$ coordinate (measured from the lower edge of the simulation box) is set to be 7.5 Mpc for the primary cluster, and 7.50, 7.96 or 8.43 Mpc for the secondary cluster, corresponding to different impact parameters. The relative velocity between the two merging clusters is fixed at 1200 km/s in the $x$ direction and 0 in the $y$ direction. The initial velocities of each cluster are set to ensure that the total momentum of the system is zero.
To demonstrate the impact of cooling and AGN feedback, we conduct three sets of merger simulations. The $m[1-9]$ simulations are mergers only, excluding both cooling and AGN feedback. The $mc[1-9]$ simulations include mergers with cooling but no AGN feedback. Our primary focus, the $mcf[1-9]$ simulations, involves mergers with both cooling and AGN feedback. All simulations are run for 10 Gyr, with data analyzed at intervals of 0.1 Gyr.
@@ 2.2 AGN Feedback
where $M_{\mathrm{cold}}$ is the total cold gas mass inside $R_{\mathrm{acc}}$, and $t_{\mathrm{ff}}=\sqrt{2r/g}$ is the free-fall time at $R_{\mathrm{acc}}$ ($r$ is the radius from the SMBH and $g$ is the gravitational acceleration). This is also known as ‘chaotic cold accretion’, which has been shown to be the dominant driver of the AGN self-regulation over the long term, both from theoretical and observational evidences (Gaspari et al. 2020, for a review). The amount of accreted gas mass is added to the SMBH mass $M_{\mathrm{BH}}$. Note that we neglect gas depletion in our official runs, but we have verified that the result is insensitive to this effect by testing one of the $mcf$ simulation.
@@
where $\dot{M}$ and $\dot{E}$ are the mass and total energy injection rates, respectively, $\eta=1$ is the mass loading factor, $\epsilon_{f}=0.01$ is the feedback efficiency, and $c$ is the speed of light. The injection is applied to a cylinder around the central SMBH with a half height of $h_{\mathrm{ej}}=$ 2.0 kpc and radius of $r_{\mathrm{ej}}=$ 2.5 kpc. These parameter values are chosen to match those adopted in previous simulations (e.g., Yang et al. 2019; see also Yang et al. 2012 for a detailed parameter study). For each simulation grid cell inside the cylinder, the fluid variables are updated as follows:
@@
The jet outflow is assumed to be bipolar with no opening angle. Following Davé et al. (2019), the jet direction is chosen to align with the averaged angular momentum within $R_{\mathrm{acc}}$. To ensure a smooth injection of energy within the jet, we modulate the amount of injected energy using a sine function (Molnar et al., 2017): $\Delta E=E_{\mathrm{norm}}\,\mathrm{sin}(\pi h/2h_{\mathrm{ej}})$, where $E_{\mathrm{norm}}$ is the normalized constant energy increment, and $h$ is the distance from the SMBH to the grid in the direction parallel to the jet. Consequently, grid cells closer to the SMBH receive less injected energy, while those further away receive more. The sine function is normalized to ensure that the total injected energy is consistent with Eq. 8. To preserve the symmetry of the jet from the perspective of the SMBH, injections are performed in the SMBH reference frame.
2.3 Black Hole Dynamics
In our simulations, a SMBH is represented by a particle at the center of each galaxy cluster, performing AGN feedback to regulate the cluster core. The initial BH masses are determined by $M_{\mathrm{tot},500}$ using the observed relation of $M_{\mathrm{BH}}-M_{\mathrm{tot},500}$ (Gaspari et al. 2019, Figure 8), where $M_{\mathrm{tot},500}$ is the cluster’s total mass within a radius $R_{500}$ that contains an average total mass density 500$\times$ the critical cosmic density. For our three clusters, the initial SMBH masses are $7.1\times 10^{10}M_{\odot}$, $1.5\times 10^{10}M_{\odot}$, $3.1\times 10^{9}M_{\odot}$, respectively.
@@ 2.3 Black Hole Dynamics
2.4 Radiative Cooling
In this work, we adopt the exact integration scheme to solve the cooling source term (Townsend, 2009; Gaspari et al., 2012). This method allows us to solve the ordinary differential equation of radiative cooling precisely, with no time-step constraints needed. We follow the same module described in Gaspari et al. (2012) and implement it into GAMER. We apply a temperature floor of $10^{4}$ K. An analytic test for the exact cooling scheme is described in Appendix B.
3 Results
In this section, we start by comparing the three sets of simulations including different physical processes (i.e., the $m$, $mc$, and $mcf$ simulations) in Section 3.1. We show that AGN feedback is essential for balancing cooling and that the $mcf$ simulations (i.e., mergers incorporating cooling and AGN feedback) are the most realistic setups. In Section 3.2, we analyze the entropy evolution of the $mcf$ simulations for different merger parameters and examine the transition from CCs to NCCs. In Section 3.3, we focus on the interplay of heating and cooling and investigate the impact of cluster merger and AGN feedback on CC destruction.
3.1 Mergers With and Without AGN Feedback
@@
Nevertheless, when AGN feedback is included, the central entropy is stabilized. The kinetic jets heat and push the cold gas in the core region outward, thereby reducing gas density and increasing temperature, effectively halting the runaway cooling process. In all of our $mcf$ simulations, AGN feedback heats the cluster cores and increases the central entropy to $\sim 30-1000$ keV cm2, consistent with the observed profiles in Cavagnolo et al. (2009). The clusters achieve self-regulation when both cooling and AGN feedback are incorporated. Our findings indicate that AGN feedback is essential for balancing cooling and preventing cooling catastrophes in CC clusters, consistently with previous works (e.g., Gaspari et al. 2011, 2012; Yang & Reynolds 2016). Therefore, realistic simulations must include both cooling and (mechanical) AGN feedback to achieve a long-term self-regulation.
@@
Figure 1: Entropy profiles of merger simulations at 10 Gyr with different physics processes, compared to the initial profile. Each shaded region corresponds to the scattering among the nine runs with varying $R$ and $b$. The $mc$ simulations lead to overcooling cluster cores, while the $mcf$ simulations result in realistic entropy profiles. Note that the low-entropy region at large radii in the initial profile is due to the secondary cluster.
3.2 Entropy Evolution for Different Merger Parameters
@@
Figure 2: Central entropy $S_{40}$ evolution for each $mcf$ (solid lines) and $m$ (dashed lines) simulation. Left: $R=1$ simulations. Center: $R=3$ simulations. Right: $R=10$ simulations. The two vertical dotted lines of the same color represent the times of first core passage and BH merger. The horizontal black dash-dotted line at $S_{40}=50~{}$keV cm2 separates CCs and NCCs. The grey solid lines ($scf$) represent the single cluster evolution for comparison.
@@
We investigate the entropy evolution of the nine $mcf$ simulations and characterize the transition from CCs to NCCs during the mergers. Figure 2 shows the central entropy evolution of each $mcf$ simulation, with three subplots each displaying three cases with the same mass ratio $R$. The simulations with the same impact parameter $b$ are plotted in the same color for comparison. A horizontal black line at $S_{40}=50~{}$keV cm2 separates CCs and NCCs. The two vertical dotted lines of the same color represent the times of the first core passage (defined as the time when the distance between the two SMBHs reaches the first minimum) and SMBH merger. The period of time between these two lines indicates the epoch when the merger is effective. Note that for the $R=10$ simulations, the SMBHs have not merged at 10 Gyr. The $m$ simulations are also presented in the dashed lines with the same color for comparison.
@@
To demonstrate the influence of cluster mergers, we include another case for comparison purposes, which is a single, isolated cluster simulation with feedback and cooling (referred to as the $scf$ simulation). In this case, there is no merger and AGN feedback self-regulates the cooling ICM. When there is strong cooling which lowers the entropy, powerful jet outbursts occur which act to raise the entropy. As a result, the system presents a quasi-periodic variation in the central entropy, as depicted by the grey line in Figure 2.
3.2.1 Mergers with Mass-Ratio $R=1$
@@
Figure 3: Entropy slices on the $z=7.5$ plane for simulation $mcf1$ ($R=1$, $b=0$). The epochs are t = 0, 1, 2, 3, 4, 5, 6, 8, 10 Gyr. Each panel is 5 Mpc on a side. The cyan dot represents the SMBH of the primary cluster, while the red dot represents the SMBH of the secondary cluster. The core remains in its CC state after merger.
@@
Figure 3 presents the 2D entropy maps on the $z=7.5$ plane (the middle of the simulation box) of the $mcf1$ simulation at various times, illustrating significant entropy changes corresponding to the $S_{40}$ evolution. During the first core passage at $t\approx 1.6$ Gyr, the initial CCs are directly heated by shocks, and the low-entropy gas in the cluster cores is stripped, forming a disk-like structure between the clusters. As the clusters reach their maximum separation at $t\approx 2.9$ Gyr, they begin to fall back toward each other, causing the gas to be dragged back and mixed further. The clusters oscillate several times before finally merging into a single cluster at $t\approx 3.8$ Gyr. After the merger, cooling dominates the system, leading to a decrease in central entropy. At later stages, the entropy maps clearly reveal the reformation of a CC structure.
Comparing $mcf1$ to the merger-only run ($m1$), the $S_{40}$ evolution of the two simulations begins to diverge after the peak of the first core passage. The second and third peaks in $mcf1$ are weaker than those in $m1$. Following the SMBH merger, $S_{40}$ in $mcf1$ consistently declines due to the inclusion of radiative cooling, whereas in $m1$, it remains at the same level until the end of the simulation. This indicates that, when cooling is included, heating from the cluster merger alone is insufficient to turn the CC into NCC in this case.
The off-axis mergers exhibit a markedly different evolution compared to the head-on merger. During the first core passage, the clusters sideswipe each other, stripping the low-entropy gas from the opposing cluster. As they fall back and rotate around each other, the gas is thoroughly mixed and heated, resulting in a significant increase in central entropy. By the time the SMBHs merge, the central entropy in the $mcf2$ and $mcf3$ simulations has reached $S_{40}\approx$ 500-600 keV cm2. Although it gradually declines over time, the central entropy remains high at the end of the simulation. Consequently, the final core remnants of these mergers are NCCs. This indicates that the equal-mass off-axis mergers result in strong heating due to gas mixing and sloshing and can destroy the CCs into NCCs, consistent with previous merger-only simulations (ZuHone et al., 2010).
We observe an interesting difference when comparing the $mcf2$ and $mcf3$ simulations with their merger-only counterparts ($m2$, $m3$). The central entropy values are 100-200 keV cm2 higher in the $mcf$ simulations than in the $m$ simulations post-merger, despite the cooling in the $mcf$ simulations. We attribute this to an overheating phenomenon caused by AGN feedback, which we discuss in detail in Section 3.3.
3.2.2 Mergers with Mass-Ratio $R=3$

### v3

@@ I Introduction
Over the past decade, simulations of CC clusters incorporating mechanical AGN feedback and radiative cooling have successfully produced self-regulated systems and CC profiles consistently with observations (Gaspari et al. 2011, 2012, Li & Bryan 2014, Prasad et al. 2015, Yang & Reynolds 2016, Yang & Reynolds 2016). On the other hand, isolated cluster simulations investigating whether AGN feedback is capable of destroying the CCs have not reached a definitive conclusion. Guo & Oh (2009) and Barai et al. (2016) find that AGN feedback can transform CCs into NCCs, while Gaspari et al. (2013) and Ehlert et al. (2022) show that an efficient self-regulation preserves the CC structure over the long term, in particular through chaotic cold accretion. In order to understand the role of AGN feedback in the CC/NCC dichotomy, it is critical to investigate the interplay between radiative cooling and AGN feedback/feeding during diverse cluster merger events.
@@
The structure of the paper is as follows. In Section II, we describe the setups of our cluster merger simulations and the modules of AGN feedback and radiative cooling. In Section III, we present the comparison of merger simulations with different physical processes in Section III.1, the entropy evolution of mergers with AGN feedback and radiative cooling in Section III.2, and the processes of heating and cooling in Section III.3. In Section IV, we compare our results with previous works in Section IV.1, discuss the implications of CC/NCC transitions in Section IV.2, and address the limitations of our simulations in Section IV.3. Finally, we summarize our findings in Section V.
II Methods
We perform 3D hydrodynamic simulations using GAMER (Schive et al., 2018), a GPU-accelerated adaptive mesh refinement (AMR) code. With GPU-acceleration, GAMER outperforms FLASH (Fryxell et al., 2000) by nearly two orders of magnitude in cluster merger simulations. Our main focus is the mergers of galaxy clusters incorporating AGN feedback and radiative cooling, but we also conduct mergers only and mergers without AGN feedback for comparison. We vary the mass ratios and impact parameters of cluster mergers to assess their respective impacts on CC destruction.
In GAMER, cold dark matter is represented by particles, which are associated with leaf patches and updated alongside them. The gravitational potential for both particles and fluid is computed by solving the discretized Poisson equation subject to given boundary conditions, utilizing the fast Fourier transform method on the root level and iterative methods on refined levels. Gravity is coupled to hydrodynamics in an operator-unsplit predictor-corrector approach, which is second-order accurate.
II.1 Simulation Setup
@@
[…]
@@
[…]
@@
We perform 3D simulations of mergers between a primary and a secondary cluster within a cubic domain of (15 Mpc)3. “Diode” boundary conditions for the hydrodynamics are adopted to permit outflow while preventing inflow. DM particles have “outflow” boundary conditions; they are lost once they exit the domain. Isolated boundary conditions are used for the gravitational potential. The cluster mergers are assumed to occur on the xyx-y plane. For each simulation, the $x$ coordinates of the two merging clusters are fixed at -1.76 Mpc and 1.76 Mpc, respectively, measured from the center of the simulation box. The $y$ coordinate is set to be 0 Mpc for the primary cluster, and 0, 0.46 or 0.93 Mpc for the secondary cluster, corresponding to different impact parameters. The relative velocity between the two merging clusters is fixed at 1200 km/s in the $x$ direction and 0 in the $y$ direction. The initial velocities of each cluster are set to ensure that the total momentum of the system is zero.
To demonstrate the impact of cooling and AGN feedback, we conduct three sets of merger simulations. The m[19]m[1-9] simulations are mergers only, excluding both cooling and AGN feedback. The mc[19]mc[1-9] simulations include mergers with cooling but no AGN feedback. Our primary focus, the mcf[19]mcf[1-9] simulations, involves mergers with both cooling and AGN feedback. All simulations are run for 10 Gyr, with data analyzed at intervals of 0.1 Gyr.
@@ II.2 AGN Feedback
where $M_{\mathrm{cold}}$ is the total cold gas mass inside $R_{\mathrm{acc}}$, and $t_{\mathrm{ff}}=\sqrt{2r/g}$ is the free-fall time at $R_{\mathrm{acc}}$ ($r$ is the radius from the SMBH and $g$ is the gravitational acceleration). This is also known as ‘chaotic cold accretion’, which has been shown to be the dominant driver of the AGN self-regulation over the long term, both from theoretical and observational evidences (Gaspari et al. 2020, for a review). In reality, the micro-accretion rate of the central SMBH can differ significantly from the Bondi rate tied to hot gas (e.g., Cho et al. 2024) and the free-fall estimate tied to cold gas (e.g., Guo et al. 2024). However, as these processes remain unresolved, we adopt the hot and cold accretion models for consistency with previous studies on AGN feedback on cluster scales. In our simulations, when cold gas is present within the accretion radius, the total accretion rate increases significantly due to the contribution from cold accretion; at other times, it is mainly dominated by Bondi accretion. In general, powerful episodes of AGN activity are triggered by the cold accretion, whereas feedback due to the hot accretion is more continuous and gentle. The amount of accreted gas mass is added to the SMBH mass $M_{\mathrm{BH}}$. Note that we neglect gas depletion in our official runs, but we have verified that the result is insensitive to this effect by testing one of the mcfmcf simulation.
@@
where $\dot{M}$ and $\dot{E}$ are the mass and total energy injection rates, respectively, $\eta=1$ is the mass loading factor, $\epsilon_{f}=0.01$ is the feedback efficiency, and $c$ is the speed of light. The injection is applied to a cylinder around the central SMBH with a half height of $h_{\mathrm{ej}}=$ 2.0 kpc and radius of $r_{\mathrm{ej}}=$ 2.5 kpc. On the micro scale of the BH accretion disk, the feedback efficiency can vary between 0.01 and 1, depending on the BH spin and the plasma properties (e.g., Tchekhovskoy et al. 2010). However, on the macro scale of galaxy clusters, the relationship between the mass injection rate and the measured central inflow rate is less clear, with feedback efficiencies adopted in previous simulations typically ranging between $\sim 10^{-3}-10^{-1}$ (Gaspari et al., 2011; Li et al., 2015; Yang & Reynolds, 2016). Here we adopt an intermediate value of $\epsilon_{f}=0.01$ as a representative case, as also supported by physical models of AGN self-regulation linking micro to macro scales (Gaspari & Sądowski 2017). Note that, while self-regulated AGN feedback could be established with feedback efficiencies in the above range, a higher feedback efficiency would tend to produce more variable AGN activity, and a lower feedback efficiency would yield more gentle fluctuations around the quasi-equilibrium level (see Yang et al. 2012 for a detailed parameter study). For each simulation grid cell inside the cylinder, the fluid variables are updated as follows:
@@
The jet outflow is assumed to be bipolar with no opening angle. Following Davé et al. (2019), the jet direction is chosen to align with the averaged angular momentum within $R_{\mathrm{acc}}$. To ensure a smooth injection of energy within the jet, we modulate the amount of injected energy using a sine function (Molnar et al., 2017): $\Delta E=E_{\mathrm{norm}}\,\mathrm{sin}(\pi h/2h_{\mathrm{ej}})$, where $E_{\mathrm{norm}}$ is the normalized constant energy increment, and $h$ is the distance from the SMBH to the grid in the direction parallel to the jet. Consequently, grid cells closer to the SMBH receive less injected energy, while those further away receive more. The sine function is normalized to ensure that the total injected energy is consistent with Eq. 8. To preserve the symmetry of the jet from the perspective of the SMBH, injections are performed in the SMBH reference frame. After computing $\Delta E$ and $\Delta m$ of each grid cell, the magnitude of momentum is calculated following Eq. 10. By keeping the momentum perpendicular to the jet axis unchanged, the momentum parallel to the jet axis is updated. The momentum is then transformed from the SMBH reference frame back to the simulation rest frame.
II.3 Black Hole Dynamics
[…]
@@ II.3 Black Hole Dynamics
II.4 Radiative Cooling
ICM radiative cooling is typically modeled as a combination of bremsstrahlung ($T\gtrsim 10^{7}$ K) and line emission ($T\sim 10^{4}-10^{7}$ K). In this work, we adopt the exact integration scheme to solve the cooling source term (Townsend, 2009; Gaspari et al., 2012). This method allows us to solve the ordinary differential equation of radiative cooling precisely, with no time-step constraints needed. We follow the same module described in Gaspari et al. (2012) and implement it into GAMER. The cooling function is modeled using an analytical fit to the Sutherland & Dopita (1993) tabulated values for a fully ionized plasma with metallicity $Z=0.3Z_{\odot}$. We apply a temperature floor of $10^{4}$ K. An analytic test for the exact cooling scheme is described in Appendix B.
III Results
In this section, we start by comparing the three sets of simulations including different physical processes (i.e., the $m$, mcmc, and mcfmcf simulations) in Section III.1. We show that AGN feedback is essential for balancing cooling and that the mcfmcf simulations (i.e., mergers incorporating cooling and AGN feedback) are the most realistic setups. In Section III.2, we analyze the entropy evolution of the mcfmcf simulations for different merger parameters and examine the transition from CCs to NCCs. In Section III.3, we focus on the interplay of heating and cooling and investigate the impact of cluster merger and AGN feedback on CC destruction.
III.1 Mergers With and Without AGN Feedback
@@
Nevertheless, when AGN feedback is included, the central entropy is stabilized. The kinetic jets heat and push the cold gas in the core region outward, thereby reducing gas density and increasing temperature, effectively halting the runaway cooling process. In all of our mcfmcf simulations, AGN feedback heats the cluster cores and increases the central entropy to $\sim 30-1000$ keV cm2, consistent with the observed profiles in Cavagnolo et al. (2009). The clusters achieve self-regulation when both cooling and AGN feedback are incorporated. Our findings indicate that AGN feedback is essential for balancing cooling and preventing cooling catastrophes in CC clusters, consistently with previous works (e.g., Gaspari et al. 2011, 2012; Yang & Reynolds 2016, 2016). Therefore, realistic simulations must include both cooling and (mechanical) AGN feedback to achieve a long-term self-regulation.
@@
Figure 1: Entropy profiles of merger simulations at 10 Gyr with different physics processes, compared to the initial profile. Each shaded region corresponds to the scattering among the nine runs with varying $R$ and $b$. The $mc$ simulations lead to overcooling cluster cores, while the mcfmcf simulations result in realistic entropy profiles. Note that the low-entropy region at large radii in the initial profile is due to the secondary cluster.
III.2 Entropy Evolution for Different Merger Parameters
@@
Figure 2: Central entropy $S_{40}$ evolution for each mcfmcf (solid lines) and $m$ (dashed lines) simulation. Left: $R=1$ simulations. Center: $R=3$ simulations. Right: $R=10$ simulations. The two vertical dotted lines of the same color represent the times of first core passage and BH merger. Note that for the $R=10$ simulations, the BHs have not merged at 10 Gyr. The horizontal black dash-dotted line at $S_{40}=50$~keV cm2 separates CCs and NCCs. The grey solid lines (scfscf) represent the single cluster evolution for comparison.
@@
We investigate the entropy evolution of the nine mcfmcf simulations and characterize the transition from CCs to NCCs during the mergers. Figure 2 shows the central entropy evolution of each mcfmcf simulation, with three subplots each displaying three cases with the same mass ratio $R$. The simulations with the same impact parameter $b$ are plotted in the same color for comparison. A horizontal black line at $S_{40}=50$~keV cm2 separates CCs and NCCs. The two vertical dotted lines of the same color represent the times of the first core passage (defined as the time when the distance between the two SMBHs reaches the first minimum) and SMBH merger. The period of time between these two lines indicates the epoch when the merger is effective. Note that for the $R=10$ simulations, the SMBHs have not merged at 10 Gyr. The $m$ simulations are also presented in the dashed lines with the same color for comparison.
@@
To demonstrate the influence of cluster mergers, we include another case for comparison purposes, which is a single, isolated cluster simulation with feedback and cooling (referred to as the scfscf simulation). In this case, there is no merger and AGN feedback self-regulates the cooling ICM. When there is strong cooling which lowers the entropy, powerful jet outbursts occur which act to raise the entropy. As a result, the system presents a quasi-periodic variation in the central entropy, as depicted by the grey line in Figure 2.
III.2.1 Mergers with Mass-Ratio $R=1$
@@
Figure 3: Entropy slices on the z=0z={0} plane for simulation $mcf1$ ($R=1$, $b=0$). The epochs are t = 0, 1, 2, 3, 4, 5, 6, 8, 10 Gyr. Each panel is 5 Mpc on a side. The cyan dot represents the SMBH of the primary cluster, while the red dot represents the SMBH of the secondary cluster. The core remains in its CC state after merger.
@@
Figure 3 presents the 2D entropy maps on the $z=0$ plane (the middle of the simulation box) of the $mcf1$ simulation at various times, illustrating significant entropy changes corresponding to the $S_{40}$ evolution. During the first core passage at $t\approx 1.6$ Gyr, the initial CCs are directly heated by shocks, and the low-entropy gas in the cluster cores is stripped, forming a disk-like structure between the clusters. As the clusters reach their maximum separation at $t\approx 2.9$ Gyr, they begin to fall back toward each other, causing the gas to be dragged back and mixed further. The clusters oscillate several times before finally merging into a single cluster at $t\approx 3.8$ Gyr. After the merger, cooling dominates the system, leading to a decrease in central entropy. At later stages, the entropy maps clearly reveal the reformation of a CC structure.
Comparing $mcf1$ to the merger-only run ($m1$) in Figure 2, the $S_{40}$ evolution of the two simulations begins to diverge after the peak of the first core passage. The second and third peaks in $mcf1$ are weaker than those in $m1$. Following the SMBH merger, $S_{40}$ in $mcf1$ consistently declines due to the inclusion of radiative cooling, whereas in $m1$, it remains at the same level until the end of the simulation. This indicates that, when cooling is included, heating from the cluster merger alone is insufficient to turn the CC into NCC in this case.
The off-axis mergers exhibit a markedly different evolution compared to the head-on merger. During the first core passage, the clusters sideswipe each other, stripping the low-entropy gas from the opposing cluster. As they fall back and rotate around each other, the gas is thoroughly mixed and heated, resulting in a significant increase in central entropy. By the time the SMBHs merge, the central entropy in the $mcf2$ and $mcf3$ simulations has reached $S_{40}\approx$ 300-500 keV cm2. Although it gradually declines over time, the central entropy remains high at the end of the simulation. Consequently, the final core remnants of these mergers are NCCs. This indicates that the equal-mass off-axis mergers result in strong heating due to gas mixing and sloshing, consistent with previous merger-only simulations (ZuHone et al., 2010). However, as will be discussed later in Section III.3, AGN heating still plays a significant role in transforming the CCs into NCCs.
III.2.2 Mergers with Mass-Ratio $R=3$